OpenStudy (anonymous):

identify the vertical asymptotes of f(x)=x+6/x^2-9x+18

OpenStudy (anonymous):

start with the meaning of an asymptote.

OpenStudy (anonymous):

i got x=6 and x=3

OpenStudy (anonymous):

asymptote- a straight line that continually approaches a given curve but does not meet it at any finite distance. here, the straight line(asymptote) is a line parallel to the y axis.

OpenStudy (anonymous):

@Abhishek619 am i right?

OpenStudy (anonymous):

i'm sorry, no. just think over it.

OpenStudy (anonymous):

http://www.purplemath.com/modules/asymtote.htm

OpenStudy (anonymous):

ok i tried again and got -6 and x=3

OpenStudy (anonymous):

@jeremy02030203 nice explanation. :) @helpineedit - go through the link mentioned above, try to solve then. :D post the answer.

OpenStudy (anonymous):

oops sorry, is your question \[\frac{ x+6 }{ x ^{2}-9x+18 }\]?

OpenStudy (amoodarya):

find the roots of x^2-9x+18

OpenStudy (anonymous):

As the link I posted says, you can find the vertical asymptotes by setting the denominator to 0. x^2-9x+18=0 which is (x-3)(x-6)=0 therefore a value or 3 or 6 will give you 0

OpenStudy (anonymous):

yes @Abhishek619

OpenStudy (anonymous):

and the answer i have is -6 and -3

OpenStudy (anonymous):

I'm sorry, I read your question wrong! My bad! The solution which you obtained at first is the answer to your question.

OpenStudy (anonymous):

yeah your first answer was right

OpenStudy (anonymous):

i got x=6 and x=3 that one @jeremy02030203

OpenStudy (anonymous):

x=3 and x=6